### v2

@@ 3.1 Extending CMP to Weighted Case
3.1.1 Vanilla CMP Mechanism.
The CMP mechanism is a simple way to combine agent reports with a predicted optimal location. It does this by taking $P^{\prime}$ as the multiset of some copies of predicted location and applying it to GCM mechanism. The number of copies is controlled by a confidence parameter $c\in[0,1)$, more specifically, it adds $cn$ copies of $\hat{o}$.
@@
In this work, we extend the Coordinate Median with Prediction (CMP) mechanism to the weighted strategic setting, and analyze its trade-off between consistency and robustness.
@@
3.1.2 Extension of CMP to Weighted Setting
We now analyze the Coordinate Median with Prediction (CMP) mechanism in the weighted setting introduced in our contribution section. As before, CMP takes as input the reported agent locations and a predicted optimal location, adds phantom points at the prediction, and returns the coordinate-wise median over the augmented set. The mechanism does not consider weights when computing the facility location, and thus remains deterministic and strategyproofness by construction.
In the weighted setting, the key difference lies in how performance is evaluated. Each agent $i$ is assigned a public weight $w_{i}>0$, and the cost function becomes:
@@ 3.1.2 Extension of CMP to Weighted Setting
Our main result, as stated in Theorem 1, shows that CMP achieves a consistency-robustness trade-off parameterized by $c\in[0,1)$, with explicit performance bounds depending on the weight ratio $W_{\max}/W_{\min}$. When all weights are equal, these bounds reduce to the known results for the unweighted case [1]. When $c=0$, CMP reduces to the original coordinate median mechanism and recovers the robustness bounds in [30].
We next analyze CMP’s behavior on a structured input family, known as the Clusters-on-Axes (COA) construction, to derive its consistency and robustness bounds and understand why these bounds are tight.
3.2 The Consistency
@@
This setup is designed to maximize the approximation ratio by concentrating high-weight agents near the optimal point and low-weight agents near the prediction. The following result formalizes that this weighting indeed leads to the worst-case consistency.
@@
This result demonstrates how the consistency of CMP degrades as the weight imbalance $\frac{W_{\max}}{W_{\min}}$ increases. In the special case where all agents have equal weights, the bound simplifies to the known result in the unweighted case. Now we begin to proof why COA can achieve this bound.
@@
Before we begin the proof, we firstly present what CA and OA is. ∎
@@
      (Less points in left)
@@
      (Less points in bottom)
@@
      (Symmetric x-clusters) If $(-x_{1},0),(x_{2},0)\in P$, then $x_{o}+x_{1}=x_{2}-x_{o}$;
@@
      (Symmetric y-clusters) If $(0,-y_{1}),(0,y_{2})\in P$, then $y_{o}+y_{1}=y_{2}-y_{o}$.
@@
3.3 The Robustness
We now analyze the robustness of CMP in the weighted setting. Robustness captures how poorly the mechanism can perform when the prediction is arbitrarily wrong. Following the same weighted COA construction as Section 3.2, we now assume that the prediction $\hat{o}$ is at $(0,0)$, simulating the wrong prediction. The CMP mechanism still outputs $(0,0)$.
@@ 3.3 The Robustness
3.4 The Trade-off Between Consistency and Robustness
@@ 3.5 Impossibility of Simultaneous Optimal Trade-off
With Corollary 3 in [23], we know that any deterministic, strategyproof, anonymous, and unanimous mechanism can be expressed as a Generalized Coordinatewise Median (GCM) mechanism with $n-1$ constant points in $P$.

### v3

@@ 3.1 Extending CMP to Weighted Case
Vanilla CMP Mechanism.
The vanilla CMP mechanism (i.e., the original version of CMP) is a simple way to combine agent reports with a predicted optimal location. It does this by taking $P^{\prime}$ as the multiset of some copies of predicted location and applying it to GCM mechanism. The number of copies is controlled by a confidence parameter $c\in[0,1)$, more specifically, it adds $cn$ copies of $\hat{o}$.
@@
In this work, we extend the Coordinate Median with Prediction (CMP) mechanism to the weighted strategic setting, and analyze its consistency and robustness.
@@
Extension of CMP to Weighted Setting.
We now analyze the Coordinate Median with Prediction (CMP) mechanism in the weighted setting introduced in our contribution section. Similar with the unweighted setting, heAs before, CMP still takes as input the reported agent locations and a predicted optimal location, adds phantom points at the prediction, and returns the coordinate-wise median over the augmented set. TThe procedure of our mechanism does not rely on the weights,consider weights when computing the facility location, and thus remains deterministic and strategyproofness by construction.
Although sharing a similar description of procedure with the vanilla CMP, the extended CMP asks for novel analysis techniques to incorporate the non-uniform weightsIn the weighted setting, the key difference lies in how performance is evaluated. In particular, eEach agent $i$ is assigned a public weight $w_{i}>0$, and the cost function becomes:
@@
Our main result, as stated in Theorem 1, shows that CMP achieves a consistency-robustness trade-off parameterized by $c\in[0,1)$, with explicit performance bounds depending on the weight ratio $W_{\max}/W_{\min}$. When theall weights are identicequal, these bounds reduce to the known results for the unweighted case [1]. Particularly, when $c$ is set to 00When $c=0$, the extended CMP reduces to the vanilla versionoriginal coordinate median mechanism and recovers the robustness bounds in [32].
We next analyze CMP’s behavior on a structured input family, known as the Clusters-on-Axes (COA) construction proposed by Agrawal et al. [1], to derive its consistency and robustness bounds and demonstrateunderstand why these bounds are tight.
3.2 Consistency Analysis
@@
This setup is designed to achieve the largest possiblmaximize the approximation ratio by concentrating high-weight agents near the optimal point and low-weight agents near the prediction. The following result formalizes that this setweighting indeed leads to the worst-case consistency.
@@
This result demonstrates how the consistency of CMP degrades as the weight imbalance $\frac{W_{\max}}{W_{\min}}$ increases. In the special case where all agents have equal weights, the bound reducsimplifies to the known result in the unweighted case. Now we begin to proveof why COA can achieve this bound.
@@
Before we begin the proof, we firstly present CA and OA proposed by Agrawal et al. [1]. ∎
@@
      Less points in left:
@@
      Less points in bottom:
@@
      Symmetric x-clusters: If $(-x_{1},0),(x_{2},0)\in P$, then $x_{o}+x_{1}=x_{2}-x_{o}$;
@@
      Symmetric y-clusters: If $(0,-y_{1}),(0,y_{2})\in P$, then $y_{o}+y_{1}=y_{2}-y_{o}$.
@@
3.3 Robustness Analysis
We now analyze the robustness of CMP in the weighted setting. Robustness captures how poorly the mechanism can perform when the prediction is arbitrarily wrong. Following the same weighted COA construction as Definition 1, we now assume that the prediction $\hat{o}$ is at $(0,0)$, simulating the wrong prediction. The CMP mechanism still outputs $(0,0)$.
@@ 3.3 Robustness Analysis
3.4 Trade-off Between Consistency and Robustness
@@ 3.5 Impossibility of Simultaneous Optimal Trade-off
Generalized Coordinate-wise Median (GCM) mechanism proposed by Procaccia et al. [25] assumes that the locations are in the (low-dimensional) Euclidean space and computes the median of the reported locations along each coordinate.
With Corollary 3 in [24], we know that any deterministic, strategyproof, anonymous, and unanimous mechanism can be expressed as a GCM mechanism with $n-1$ constant points in $P$.